OpenStudy (anonymous):

find the tangent plane to the surface of revolution obtained by rotating y = e^x about the x axis at the point (0, sqrt(2)/2,sqrt(2)/2)

OpenStudy (anonymous):

so ive got the parametrization of the surface as \[<x, e^xcos(\theta),e^xsin(\theta)>\]

OpenStudy (anonymous):

With the above i get x = 0 and theta = pi/4

OpenStudy (anonymous):

i get two vectors as \[<0,\cos(\theta),\sin(\theta)>\] and \[<x, \frac{\sqrt{2}}{2}e^x, \frac{\sqrt{2}}{2}e^x>\]

OpenStudy (anonymous):

derivative of first vector \[<0,-\sin(\theta),\cos(\theta)>\] and second one \[<1, \frac{\sqrt{2}}{2}e^x, \frac{\sqrt{2}}{2}e^x>\]

OpenStudy (anonymous):

evaluate first vector at pi/4\[<0, - \frac{\sqrt{2}}{2}, \frac{\sqrt{2}}{2}>\] and second one a x = 0\[<1, \frac{\sqrt{2}}{2}, \frac{\sqrt{2}}{2}>\]

OpenStudy (anonymous):

take cross product and get \[-i - \frac{\sqrt{2}}{2}j + \frac{\sqrt{2}}{2}k\]

OpenStudy (anonymous):

Equation of the plane is \[-x - \frac{\sqrt{2}}{2}y + \frac{\sqrt{2}}{2}z = 0\]

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

@SolomonZelman @texaschic101 ?

OpenStudy (anonymous):

what do you think @dumbcow ?

OpenStudy (anonymous):

@Directrix ?

OpenStudy (dumbcow):

ok sorry i wanted to make sure i was doing it right. you did cross-product correctly but something else is off, why did you use vector <0, cos, sin> ? the tangent plane is defined using the gradient (derivative) \[\delta_x (x - x_0) + \delta_y(y - y_0) + \delta_z(z - z_0) = 0\] for point (x0,y0,z0) derivative gives : \[<1, \frac{\sqrt{2}}{2}, \frac{\sqrt{2}}{2}>\] thus equation is \[x + \frac{\sqrt{2}}{2}(y-\frac{\sqrt{2}}{2}) + \frac{\sqrt{2}}{2} (z - \frac{\sqrt{2}}{2}) = 0\]

OpenStudy (anonymous):

i got the two vectors by letting one variable be constant and letting the other variable vary so with that vector i let x be constant and theta vary which gives that vector.

OpenStudy (anonymous):

and the opposite for the other vector, that way i can cross product the two and get a normal vector.

OpenStudy (irishboy123):

the locus \[y^2 + z^2 = e ^{2x}\] \[ø = y^2 + z^2 - e ^{2x}\] \[grad (ø) = < -2 e ^{2x}, 2y, 2z > = n ^{->}\] \[n ^{->} (0, \sqrt{2}/2, \sqrt{2}/2) = <-2, \sqrt{2}, \sqrt{2} > \] \[π: - 2x + \sqrt{2}y + \sqrt{2} z = K\] \[K = 0 + \sqrt{2} • \sqrt{2}/2 + \sqrt{2} • \sqrt{2}/2 = 2\] \[π: -\sqrt 2x + y + z = \sqrt 2\]